OpenStudy (anonymous):

Find the number of digits in 2^120x125x88x5^120 . It is rather difficult

OpenStudy (anonymous):

@mathslover @.Sam.

OpenStudy (anonymous):

Is it 125 ?

OpenStudy (kc_kennylau):

Yes

OpenStudy (phi):

I would do this \[ 2^{120}\cdot 125 \cdot 88 \cdot 5^{120} \] notice that we can write this as \[ 2^{120}\cdot 5^{120} \cdot 5^3 \cdot 2^3 \cdot 11 \\ (2\cdot 5)^{120} \cdot (5 \cdot 2)^3 \cdot 11 \\ 2^{120} \cdot 10^3 \cdot 11 \]

OpenStudy (phi):

* \[ 10^{120} \cdot 10^3 \cdot 11 \]

OpenStudy (phi):

or \[ 11 \cdot 10^{123} \] to figure out how many digits that is, look for a pattern \( 11 \cdot 10^1\) = 110 or 3 digits \( 11 \cdot 10^2\) = 1100 or 4 digits the rule appears to be \( 11 \cdot 10^n\) has 2+n digits